OpenStudy (anonymous):

Two trains started from the same place at the same time and traveled in opposite directions at rates which is differed by 20mph in 5 hours they were 500 miles apart find the rate of each plane

OpenStudy (anonymous):

5x+5(x+20)=500 5x+5x+100=500 10x=400 x=40 so speeds are 40 mph and 60 mph

OpenStudy (cwrw238):

speeds are v and v + 20 v = distance / time v + v + 20 = 500 / 5 = 100 2v = 80 v = 40 speeds of the trains are 40 and 60 mph

OpenStudy (cwrw238):

thanx

OpenStudy (hba):

WELCOME TO OPEN STUDY

OpenStudy (anonymous):

oh thank you so muchhh for answering..!^^ and thanks for welcome haha.